OpenStudy (anonymous):

A Lab Manager is trying to find the number of defected items produced in two different shifts: Day and Night Shifts. Day Shift = 20 items defected out of 2000 items. Night Shift = 15 items defected out of 1300 items. Let the defectiveness of the item follow poisson distribution. If i am given a random sample group of 10 items(All are only from either Night of Day Shift) and 2 are defected out of them, find the most likely hood ratio and also find both Type 1 and Type 2 errors

OpenStudy (anonymous):

please please help me

OpenStudy (amistre64):

sorry, im not competent in the Poisson stuff. My elementary stats class didnt even mention it.

OpenStudy (anonymous):

:( Thanks.

OpenStudy (dumbcow):

hmm well to evaluate lambda you need to take each shift and scale it to 10 items 2000/200 = 10 so 20/200 = 1/10 for every 10 items on avg 1/10 are defective 1300/130 = 10 so 15/130 = 3/26 for every 10 items on avg 3/26 are defective these will be the respective lambdas then you test for probability of random sample being from either Day shift or Night shift

OpenStudy (dumbcow):

\[P(X=2) = \frac{\lambda ^{2}e^{-\lambda}}{2!}\]

OpenStudy (anonymous):

So is that how i find the most likely hood ratio. ... ?... How do i find the Both Type1 and Type 2 Errorrs?